OpenStudy (anonymous):

Find the distance for the following points A ( -2, -3) B (-7, -8). If necessary round your answer to the nearest tenth.

OpenStudy (ajprincess):

For two points \[(x_1,y_1)\space(x_2,y_2)\] \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] Can u do nw?

OpenStudy (anonymous):

I dont understand answer choices are 1) 11 2) 7.1 3) 7 4) -7.1

OpenStudy (anonymous):

I mean is it 7.1 after i round it

OpenStudy (ajprincess):

U hav points (-2,-3) and (-7,-8) Applying this in \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] \[=\sqrt{(-7-(-2))^2+(-8-(-3))^2}\] \[=\sqrt{(-7+2)^2+(-8+3)^2}\] \[=\sqrt{(-5)^2+(-5)^2}\] \[=\sqrt{25+25}\] \[=\sqrt{50}\] \[=7.07\] \[=7.1\]

OpenStudy (ajprincess):

is it clear nw?

OpenStudy (anonymous):

Yes, thank you

OpenStudy (ajprincess):

yw.:)